Calculus II, Fall 2014 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 13

1. Section 15.4: Problems 11, 21, 23, 26, 41, 42, 43.

2. Section 15.5: Problems 21, 35, 41, 43, 47.

3. Section 15.7: Problems 9, 11, 13, 14, 15, 19.